what is the relation between wavelength and refractive index?

is it

U1 * A1 = U2 * A2

Un = refractive index
An=wavelength in the medium n

Grade:12

5 Answers

sudhanshu kumar
12 years ago

u1a1=u1a1=u2a2=........   i.e ua=constant

srinivasa rao mekapothula
12 years ago

yes ,you r right.

the purpose of studying about refractive index (u)  is to know whether the given medium is  rarer or denser . this(u) again represents whether speed of light is more or less in that medium. speed of  light is less in a denser medium(more u value).i.e., u and speed of light are inversely proportional .

but, speed, v = f n   ;

  u  =  k/ v

u  =  k / f n .

u n = k/f = constant. f= frequency of light = constant.

hence , u1 n1 = u2 n2


Tholkappian Dakshinamoorthy
7 years ago
Wavelength is inversely proportional to refractive index .because least wavelength deflect more in a medium for red it is 1.509 and violet it is 1.52 for crown glass.
Jeevesh
7 years ago
Refractive index = A + B/(wavelength)^2Where A and B are constant.Lesser the wavelength greater the refractive index.
Jeevesh
7 years ago
Refractive index = A + B/(wavelength)^2Where A and B are constant.Lesser the wavelength greater the refractive index